OpenStudy (anonymous):

3x + y = 4 2x + y = 4 (Points : 4) (1, 7) (0, 4) (3, 12) (3, 13) Answer now please???

OpenStudy (anonymous):

This one too please i would aprisiat it a lot. 2. Use the graph to state the solution for the system. x + 2y = 4 (line a) 3x 2y = 4 (line b) (Points : 4) (2, 1) (2, 1) (2, 1) (2, 1)

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

really I have no freakin clue what you r doin I'm just finishing lines and intercepts

OpenStudy (anonymous):

The answer to the first one is b but i don't know for the second one

OpenStudy (anonymous):

thanks 4 the help

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

did u use linear combination???

OpenStudy (anonymous):

http://www.freemathhelp.com/ enter your problems in that CCA?

OpenStudy (anonymous):

what does it do?

OpenStudy (anonymous):

it calculates the problem and gives you the answer, given that you write your problem in correctly

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

freak love you too and your welcome for trying

OpenStudy (anonymous):

lol thanx JD I got the second aswer too

OpenStudy (anonymous):

congrats

OpenStudy (anonymous):

this would b one solution right?? . How many solutions are there to the following system of equations? 4x 8y = 0 3x + 8y = 7 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

would it be one solution or 2???

OpenStudy (anonymous):

3x + y = 4 2x + y = 4 answer is (0,4)

OpenStudy (anonymous):

do u kno if the 3rd one is one or 2 solutions?

OpenStudy (anonymous):

btw thanx

OpenStudy (anonymous):

what are they

OpenStudy (anonymous):

linear combination for the 3r done and that is the next one i need.

OpenStudy (anonymous):

i think that one might be 2 cause u get the answer for x and y so wouldnt that b 2?

OpenStudy (anonymous):

yes.. but i dont understand the other 2 ? yu ask

OpenStudy (anonymous):

i got help with those so that is ok i appreciate the help[ with that one tho.. I might have more give me just one second.

OpenStudy (anonymous):

oh ok and yu welcome

OpenStudy (anonymous):

can u help with this one? Solve the following system of equations. d + e = 6 d e = 4 (Points : 4) (3, 1) (5, 1) no solution infinite number of solutions

OpenStudy (anonymous):

i think its no solution

OpenStudy (anonymous):

ok thanxs i might have a couple more lol hold on

OpenStudy (anonymous):

Actually u helped me with all that u could thanx I really do appreciate the help and if i could give u more than one medal i would.

OpenStudy (anonymous):

lol you welcome.. i like helping

OpenStudy (anonymous):

6. Each child ticket for a ride costs $2, while each adult ticket costs $6. If the ride collected a total of $148, and 38 tickets were sold, how many of each type of ticket was sold? (Points : 4) 20 children and 18 adults 19 children and 19 adults 14 children and 20 adults 148 adults and 38 children Can any one help please?

OpenStudy (anonymous):

20 children and 18 adults

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

20*$2 18*$6 $40 + $108 = $148

OpenStudy (anonymous):

In a recent golf match, Tiger's score was 4 less than Phil's score. Their combined scores totaled 140. Which of the following pairs of equations can be used to determine their scores? Let p represent Phil's score and t represent Tiger's score. (Points : 4) p + 140 = t p = 4 p + t = 140 p + t = 4 p + t = 140 p t = 4 p t = 140 p + t = 4 What about this one???

OpenStudy (anonymous):

Btw thanx didierdre

OpenStudy (anonymous):

The answer to 1. is (0, 4)

OpenStudy (anonymous):

i got all of those i need help on the most resent one ^

OpenStudy (anonymous):

Okay I try

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

do u kno this one ^

OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y + 5 (Points : 4) x + y = y + 5 (y + 5) + y = 6 x + (y + 5) = 6 x = (x 6) + 5 wb this one to?????

OpenStudy (anonymous):

I got the other one i need help[ on this one ^

OpenStudy (anonymous):

It would be p + t = 140 p+t = 4

OpenStudy (anonymous):

ok how would i graph this? and thank u

OpenStudy (anonymous):

x + 2y = 2 4x y = 4

OpenStudy (anonymous):

graph that ^

OpenStudy (anonymous):

For the recent one it would be (y + 5) + y = 6

OpenStudy (anonymous):

kk do u know how to graph this x + 2y = 2 4x y = 4

OpenStudy (anonymous):

No im not sure how to grpah that. Sorry

OpenStudy (anonymous):

it ok thanx for all of ur help btw r u sure on all of ur answers. I just want to make sure I got them right???

OpenStudy (anonymous):

Yes im sure.. And your very welcome =)

OpenStudy (anonymous):

i just submited the tesat and I got a c on it I might need help holdon 1 second!

OpenStudy (anonymous):

that si all thanks for ur help.

OpenStudy (anonymous):

You can send them to me my inbox if you need more help